which is the lenth the hypotinuse of a right triangle if the lenth of the triangle is 6 and 8?
Answer by Theo(13342)

I would say 10.

There's a 3:4:5 triangle.

6:8:10 divided by 2 would be 3:4:5

3^2 + 4^2 = 5^2 becomes 9 + 16 = 25 which is right.

6^2 + 8^2 = 10^2 becomes 36 + 64 = 100 which is right.

That's because in a right triangle, a^2 + b^2 = c^2 where a and b are the sides of the triangle and c is the hypotenuse.

In your triangle, 6 and 8 are the sides a and b, while 10 is the hypotenuse c.